# Understanding Fine-tuning in Approximate Unlearning: A Theoretical Perspective

### v2

@@ Abstract
Machine Unlearning has emerged as a significant area of research, focusing on ‘removing’ specific subsets of data from a trained model. Fine-tuning (FT) methods have become one of the fundamental approaches for approximating unlearning, as they effectively retain model performance. However, it is consistently observed that naive FT methods struggle to forget the targeted data. In this paper, we present the first theoretical analysis of FT methods for machine unlearning within a linear regression framework, providing a deeper exploration of this phenomenon. Our analysis reveals that while FT models can achieve zero remaining loss, they fail to forget the forgetting data, as the pretrained model retains its influence and the fine-tuning process does not adequately mitigate it. To address this, we propose a novel Retention-Based Masking (RBM) strategy that constructs a weight saliency map based on the remaining dataset, unlike existing methods that focus on the forgetting dataset. Our theoretical analysis demonstrates that RBM not only significantly improves unlearning accuracy (UA) but also ensures higher retaining accuracy (RA) by preserving overlapping features shared between the forgetting and remaining datasets. Experiments on synthetic and real-world datasets validate our theoretical insights, showing that RBM outperforms existing masking approaches in balancing UA, RA, and disparity metrics.
@@ 1 Introduction
Machine Unlearning has emerged as a prominent area that focuses on protecting individual privacy during the model training process, particularly adhering to legislation such as ‘the right to be forgotten’ [21] under the General Data Protection Regulation (GDPR) [13]. That is, it removes certain training samples from the trained model upon their users’ data deletion request. A natural approach to machine unlearning is to retrain the model from scratch, excluding the data that needs to be forgotten; this is known as exact unlearning. However, this method is highly computationally inefficient. To address this challenge, previous research has proposed a more relaxed definition of machine unlearning, where the unlearned model only needs to be approximately similar to one retrained from scratch. This led to the development of approximate unlearning methods, such as Fine-Tuning [29, 8], Gradient Ascent [10, 27], Fisher Forgetting [1, 8], and Influence Unlearning[14].
Fine-tuning, as one of the most widely used approaches in approximate unlearning, has demonstrated its empirical effectiveness. However, it can be observed in many studies [16, 29, 8, 17, 23] and our investigations in Table 1 that while fine-tuning may maintain the utility of the model on remaining data, it struggles to forget the targeted data. This raises a natural question:
@@
Table 1: Cifar-10 Class-wise Forgetting Performance Comparing Retrain and Naive FT (Fine-Tuning) Method. The table compares Retrain and FT on CIFAR-10 across multiple evaluation metrics: Unlearning Accuracy (UA), Retaining Accuracy (RA), MIA-Efficacy, Test Accuracy (TA), Avg. Disparity, and Run-Time. Values in brackets indicate the gap between FT and the golden model (i.e., Retrain). Further explanations are provided in Section 5.2.
@@
Retrain $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$ $94.81_{\pm 0.09}$ $0.00$ $82.00$
FT $91.648.36_{\pm 3.03}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{91.64}})$ $40.76_{\pm 8.03}(59.24)$ $99.92_{\pm 0.03}(0.08)$ $94.41_{\pm 0.29}(0.40)$ $37.84$ $2.53$
  • Theoretical Understanding of Fine-Tuning. We provide the first theoretical analysis of FT methods in the context of machine unlearning within a linear regression framework. Specifically, $\mathbf{1)}$ Based on the assumption of distinct features (1), our theoretical observations, which align with empirical studies, show that the remaining loss for the fine-tuning model is zero, matching that of the golden model. Moreover, the loss of the fine-tuning model on the forgetting dataset consistently remains zero, diverging from the performance of the golden model. $\mathbf{2)}$ we extend our analysis to a more complex case when the dataset retained for model retraining shares overlapping features with the forgetting dataset. This challenges assumptions of distinct feature sets across datasets, yet the previous conclusions remain valid in this case. More discussion refers to Section 3.
  • Understanding the Benefits of Masking in Fine-Tuning. Our analysis shows that naive fine-tuning (FT) methods fail to unlearn the forgetting data because the pretrained model retains information about this data, and the fine-tuning process does not effectively alter that retention. To address this, we propose removing the forgetting component to mitigate its retention in the pre-trained model, an approach that aligns with the masking concept proposed in existing work [6], which directly masks the forgetting data. However, one critical case is omitted: when the remaining data and forgetting data share similar features, it becomes unclear whether those shared features should be preserved. In Section 4, our work proves: $\mathbf{1)}$ Masking on the pretrained model can significantly improve unlearning accuracy while preserving the retaining accuracy. $\mathbf{2)}$ When considering overlapping features, retaining them does not substantially affect unlearning accuracy, but discarding them compromises the retaining accuracy.
@@ 1.1 Related Work
Machine Unlearning Methods. [2] first defined “Unlearning” as the removal of a sample that produces the same output on the dataset as if the sample had never been trained. The natural way to solve the problem is to retrain a model from scratch in response to each data deletion request. However, retraining is not feasible due to the limited time and constrained resources. [7] provided a relaxed definition inspired by Differential Privacy [5], which only requires the unlearned model to produce results similar to those of retrain-from-scratch models. This led to the development of “approximate unlearning” methods, offering more efficient computational designs for machine unlearning. [11, 14, 19, 28, 22] provide theoretical error guarantees by focusing on the empirical risk minimization problem under this probabilistic notion of unlearning. [8] proposed an information-based procedure to remove knowledge from the trained weights, without access to the original training data. Further, [9] approximated the weights inspired by NTK theory, addressing situations where the Hessian is not informative about where the model will converge into a null space. [18] avoid the computation of Hessian by introducing a method only computing conditional independence, which identifies the Markov Blanket of parameters requiring updates. [27] proposed a regularizer to reduce the ‘verification error,’ which represents the distance between the unlearned model and a retrained-from-scratch model. [16] bears a novel teacher-student formulation to achieve better performance towards unbounded unlearning problems. [17] considers model sparsity by pruning weights before the unlearning process, thereby introducing a new unlearning paradigm. [24] incorporates the variational inference and contrastive learning approaches to address the lack of supervision information (label-agnostic).
@@
Machine Unlearning Theory. For approximate unlearning, [19, 27] explored algorithms for empirical risk minimization objectives, while [22] studied population risk minimization problems, providing theoretical guarantees on both the effectiveness of unlearning and the privacy of the data subjects. [11, 31] provided the certified radius with respect to data changes before and after removals, as well as the certified budget for data removals. For exact unlearning, [28] introduced the notion of algorithmic stability, called Total Variation (TV) stability, which is suited for achieving exact unlearning. This concept was further extended to the federated setting by [3, 26]. However, existing theoretical work has primarily focused on utility guarantees, with limited analysis explaining the successes and failures of fine-tuning methods.
@@ 2 Machine Unlearning in Linear Models
Let $D=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$ be a training dataset consisting of $n$ data points, where $\mathbf{x}_{i}$ represents the feature vector, and $y_{i}$ is the response variable for each data point in the dataset $D$. Assume that each pair $(\mathbf{x}_{i},y_{i})$ is a realization of the linear regression model: $y=\mathbf{x}^{\top}\mathbf{w}_{*}$, with $\mathbf{w}_{*}\in\mathbb{R}^{d}$ being the optimal model parameter in the overparameterized regime (${n}\ll{d}$). Machine Unlearning aims to remove (or scrub) the influence of specific training data from a trained machine learning (ML) model. Let $D_{f}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n_{f}}\subseteq D$ represents a subset whose influence we want to scrub, termed the forgetting dataset. Accordingly, the complement of $D_{f}$, termed the remaining dataset, is $D_{r}=\{(\mathbf{x}_{i},y_{i})\}_{i=n_{f}+1}^{n}=D\backslash D_{f}$. The forgetting data can be represented by stacking the feature vectors and response variables as follows:
@@
[…]
Our goal is to evaluate how the fine-tuning solution $\mathbf{w}_{t}$ differs from the golden model solution $\mathbf{w}_{g}$ which refers to retraining the model parameters from scratch over the remaining dataset $D_{r}$. Existing work has assessed machine unlearning performance from various perspectives [10, 1, 8, 25]. In this paper, we focus particularly on the Unlearning Loss (UL) and Remaining Loss (RL), which refers to the model performance on the forgetting and remaining dataset respectively. These losses are defined as:
[…]
@@ 3 Naive Fine-Tuning Methods Fail To Unlearn
In empirical studies [16, 29, 8] and Table 1, it can be observed that fine-tuning may retain the utility of a model but struggles to forget. In this section, we revisit this phenomenon in a simplified setting, aiming to gain a basic understanding of why the vanilla fine-tuning method succeeds in retaining the model’s utility on the remaining dataset but fails to forget the targeted data it was trained on.
@@ 3.1 Distinct Features
The assumption implies that each of these datasets contains features that are unique to each dataset–there is no overlap in the features present in $\mathbf{X}_{f}$ and $\mathbf{X}_{r}$. The construction of $\mathbf{X}_{f}$ and $\mathbf{X}_{r}$ can be achieved by rearranging the samples in the matrix to group non-zero features into distinct blocks. Without loss of generality, we assume a structure where each matrix contains only one zero block to clear the analysis. It can be obtained immediately from 1 that $\mathbf{w}_{*}=\mathbf{w}_{*}^{f}+\mathbf{w}_{*}^{r}$, where $\mathbf{w}_{*}^{f}$ and $\mathbf{w}_{*}^{r}$ are the optimal solution such that $\mathbf{y}^{f}=\mathbf{X}_{f}^{\top}\mathbf{w}_{*}^{f}$ and $\mathbf{y}^{r}=\mathbf{X}_{r}^{\top}\mathbf{w}_{*}^{r}$. In an ideal scenario for classification tasks, each class possesses its own unique set of features that distinctly differentiates it from other classes. We later extended our analysis to overlapping features in Section 3.2.
Theorem 1.
Suppose a model is trained by the procedure 2 and 3 separately. Under the 1, it holds that
@@
Theorem 1 presents two interesting observations: 1) The fine-tuning model can perform perfectly on the remaining dataset, which indicates that the information from training data has been preserved from the original model, $\mathbf{w}_{o}$, to the unlearned model via fine-tuning, $\mathbf{w}_{t}$. 2) The loss of the fine-tuning model on the forgetting dataset consistently remains zero, which diverges from the performance of the golden model. This suggests that the fine-tuning model is unable to forget the information it previously acquired from $\mathbf{w}_{o}$, which may be contradicted by catastrophic forgetting in continual learning [20, 4].
To illustrate the behavior of fine-tuning during the unlearning process more clearly, we consider the projective nature of learning. Firstly, the solution of Equation 2 can be represented as
@@
where $\mathbf{P}_{t}$ is the projection space of $\mathbf{X}_{t}$, the $\mathbf{I}-\mathbf{P}_{t}$ is the corresponding orthogonal space, and the $\mathbf{w}_{o}$ can be also represented $\mathbf{w}_{o}=\mathbf{P}\mathbf{w}_{*}$ with $\mathbf{P}$ being the projection space of $\mathbf{X}$. According to the property of projection Corollary 1, multiplying any data matrix by a projection matrix preserves the components of the data that lie within the subspace defined by the projection. Moreover, under the distinct features assumption 1, it holds that
@@
Therefore, the unlearned model $\mathbf{w}_{t}$ from Equation 2 decomposed into two components for the unlearning process: the first part, $\mathbf{w}_{*}^{r}$, preserves the accuracy on the remaining data, while the second part, $\mathbf{w}_{*}^{f}$, also ensures accuracy on the forget data. However, the projection of $\mathbf{w}_{*}^{f}$ onto the fine-tuning space $\mathbf{P}_{t}$ has no effect, ultimately resulting in the unlearned model $\mathbf{w}_{t}$ being exactly the same as the pretrained model $\mathbf{w}_{o}$. The proof of Theorem 1 is provided in Section B.2.
Figure 1: Machine Unlearning Performance via (Masked) Fine-tuning with (without) Overlapping Features. Figure 1(a) and Figure 1(b) present the relationship between machine unlearning loss (i.e. RA, UA) and the number of fine-tuning data samples under distinct features and overlapping features assumptions, using naive FT method. In contrast, Figure 1(c) and Figure 1(d) show the same relationship using masked fine-tuning methods, as discussed in Section 4.

### v3

@@ Abstract
Machine Unlearning has emerged as a significant area of research, focusing on ‘removing’ specific subsets of data from a trained model. Fine-tuning (FT) methods have become one of the fundamental approaches for approximating unlearning, as they effectively retain model performance. However, it is consistently observed that naive FT methods struggle to forget the targeted data. In this paper, we present a theoretical analysis—the first to characterize FT unlearning behavior in linear models, providing a deeper exploration of this phenomenon. Our analysis reveals that while FT models can achieve zero remaining loss, they fail to forget the forgetting data, as the pretrained model retains its influence and the fine-tuning process does not adequately mitigate it. To address this, we propose a novel Retention-Based Masking (RBM) strategy that constructs a weight saliency map based on the remaining dataset, unlike existing methods that focus on the forgetting dataset. Our theoretical analysis demonstrates that RBM not only significantly improves unlearning accuracy (UA) but also ensures higher retaining accuracy (RA) by preserving overlapping features shared between the forgetting and remaining datasets. Experiments on synthetic and real-world datasets validate our theoretical insights, showing that RBM outperforms existing masking approaches in balancing UA, RA, and disparity metrics.
@@ 1 Introduction
Machine Unlearning has emerged as a prominent area that focuses on protecting individual privacy during the model training process, particularly adhering to legislation such as ‘the right to be forgotten’ (rosen2011right) under the General Data Protection Regulation (GDPR) (hoofnagle2019european). That is, it removes certain training samples from the trained model upon their users’ data deletion request. A natural approach to machine unlearning is to retrain the model from scratch, excluding the data that needs to be forgotten; this is known as exact unlearning. However, this method is highly computationally inefficient. To address this challenge, previous research has proposed a more relaxed definition of machine unlearning, where the unlearned model only needs to be approximately similar to one retrained from scratch. This led to the development of approximate unlearning methods, such as Fine-Tuning (warnecke2021machine; golatkar2020eternal), Gradient Ascent (graves2021amnesiac; thudi2022unrolling), Fisher Forgetting (becker2022evaluating; golatkar2020eternal), and Influence Unlearning(izzo2021approximate). Fine-tuning, as one of the most widely used approaches in approximate unlearning, has demonstrated its empirical effectiveness. However, it can be observed in many studies (kurmanji2024towards; warnecke2021machine; golatkar2020eternal; liu2024model; sharma2024discriminative) and our investigations in Table˜1 that while fine-tuning may maintain the utility of the model on remaining data, it struggles to forget the targeted data. This raises a natural question:
@@
Table 1: Cifar-10 Class-wise Forgetting Performance Comparing Retrain and Naive FT (Fine-Tuning) Method. The table compares Retrain and FT on CIFAR-10 across multiple evaluation metrics: Unlearning Accuracy (UA), Retaining Accuracy (RA), MIA-Efficacy, Test Accuracy (TA), Avg. Disparity, and Run-Time. Values in brackets indicate the gap between FT and the golden model (i.e., Retrain). Further explanations are provided in Section˜5.2.
@@
Retrain $_{\pm 0.00} 100.00±0.0010$$_{\pm 0.00}$ $100.00_{\pm 0.00}$ $94.81_{\pm 0.09}$ $0.00$ $82.00$
[…]
  • Theoretical Understanding of Fine-Tuning. We provide a theoretical analysis—the first to characterize FT unlearning behavior in linear models. Specifically, $\mathbf{1)}$ Based on the assumption of distinct features (˜3.1), our theoretical observations, which align with empirical studies, show that the remaining loss for the fine-tuning model is zero, matching that of the golden model. Moreover, the loss of the fine-tuning model on the forgetting dataset consistently remains zero, diverging from the performance of the golden model. $\mathbf{2)}$ We extend our analysis to a more complex case when the dataset retained for model retraining shares overlapping features with the forgetting dataset. This challenges assumptions of distinct feature sets across datasets, yet the previous conclusions remain valid in this case. More discussion refers to Section˜3.
  • Understanding the Benefits of Masking in Fine-Tuning. Our analysis shows that naive fine-tuning (FT) methods fail to unlearn the forgetting data because the pretrained model retains information about this data, and the fine-tuning process does not effectively alter that retention. To address this, we propose removing the forgetting component to mitigate its retention in the pre-trained model, an approach that aligns with the masking concept proposed in existing work fan2023salun, which directly masks the forgetting data. However, one critical case is omitted: when the remaining data and forgetting data share similar features, it becomes unclear whether those shared features should be preserved. In Section˜4, our work proves: $\mathbf{1)}$ Masking on the pretrained model can significantly improve unlearning accuracy while preserving the retaining accuracy. $\mathbf{2)}$ When considering overlapping features, retaining them does not substantially affect unlearning accuracy, but discarding them compromises the retaining accuracy.
@@ 1.1 Related Work
Machine Unlearning Methods. cao2015towards first defined “Unlearning” as the removal of a sample that produces the same output on the dataset as if the sample had never been trained. The natural way to solve the problem is to retrain a model from scratch in response to each data deletion request. However, retraining is not feasible due to the limited time and constrained resources. ginart2019making provided a relaxed definition inspired by Differential Privacy (dwork2014algorithmic), which only requires the unlearned model to produce results similar to those of retrain-from-scratch models. This led to the development of “approximate unlearning” methods, offering more efficient computational designs for machine unlearning. guo2019certified; izzo2021approximate; neel2021descent; ullah2021machine; sekhari2021remember provide theoretical error guarantees by focusing on the empirical risk minimization problem under this probabilistic notion of unlearning. golatkar2020eternal proposed an information-based procedure to remove knowledge from the trained weights, without access to the original training data. Further, golatkar2020forgetting approximated the weights inspired by NTK theory, addressing situations where the Hessian is not informative about where the model will converge into a null space. mehta2022deep avoid the computation of Hessian by introducing a method only computing conditional independence, which identifies the Markov Blanket of parameters requiring updates. thudi2022unrolling proposed a regularizer to reduce the ‘verification error,’ which represents the distance between the unlearned model and a retrained-from-scratch model. kurmanji2024towards bears a novel teacher-student formulation to achieve better performance towards unbounded unlearning problems. liu2024model considers model sparsity by pruning weights before the unlearning process, thereby introducing a new unlearning paradigm. shen2024label incorporates the variational inference and contrastive learning approaches to address the lack of supervision information (label-agnostic). torkzadehmahani2024improved leverages memorization dynamics to design localized masks that better preserve task-relevant representations, while foster2024fast computes a parameter-importance ratio between the forget and retain sets using the Fisher information matrix. In contrast, our method derives the masking strategy directly from the remaining dataset, guided by theoretical insights into feature overlap and retention.
@@
Machine Unlearning Theory. For approximate unlearning, neel2021descent; thudi2022unrolling explored algorithms for empirical risk minimization objectives, while sekhari2021remember studied population risk minimization problems, providing theoretical guarantees on both the effectiveness of unlearning and the privacy of the data subjects. guo2019certified; zhang2022prompt provided the certified radius with respect to data changes before and after removals, as well as the certified budget for data removals. For exact unlearning, ullah2021machine introduced the notion of algorithmic stability, called Total Variation (TV) stability, which is suited for achieving exact unlearning. This concept was further extended to the federated setting by che2023fast; tao2024communication. chien2024langevin introduced Langevin Unlearning, which interprets noisy gradient descent as an implicit Bayesian sampling mechanism that promotes forgetting through stochastic noise injection. However, existing theoretical work has primarily focused on utility guarantees or privacy-driven dynamics, with limited analysis explaining the successes and failures of fine-tuning methods.
@@ 2 Machine Unlearning in Linear Models
[…]
@@
$\text{OT:}\quad\mathbf{w}_{o}$ $=\underset{\mathbf{w}}{\operatorname{argmin}}{\|\mathbf{w}\|_{2}},\quad\text{ s.t. }\mathbf{y}=\mathbf{X}^{\top}\mathbf{w}$ (1)
$\text{FT:}\quad\mathbf{w}_{t}$ $=\underset{\mathbf{w}}{\operatorname{argmin}}{\|\mathbf{w}-\mathbf{w}_{o}\|_{2}},\quad\text{ s.t. }\mathbf{y}_{t}=\mathbf{X}_{t}^{\top}\mathbf{w}$ (2)
$\text{GU:}\quad\mathbf{w}_{g}$ $=\underset{\mathbf{w}}{\operatorname{argmin}}{\|\mathbf{w}\|_{2}},\text{ s.t. }\mathbf{y}_{r}=\mathbf{X}_{r}^{\top}\mathbf{w}$ (3)
Our goal is to evaluate how the fine-tuning solution $\mathbf{w}_{t}$ differs from the golden model solution $\mathbf{w}_{g}$ which refers to retraining the model parameters from scratch over the remaining dataset $D_{r}$. Existing work has assessed machine unlearning performance from various perspectives (graves2021amnesiac; becker2022evaluating; golatkar2020eternal; song2019privacy). In this paper, we focus particularly on the Unlearning Loss (UL) and Remaining Loss (RL), which refers to the model performance on the forgetting and remaining dataset respectively. These losses are defined as:
$\text{RL:}\quad L(\mathbf{w},D_{r})=\frac{1}{n_{r}}\|\mathbf{X}_{r}^{\top}\mathbf{w}-\mathbf{y}_{r}\|_{2}^{2},\quad\text{UL:}\quad L(\mathbf{w},D_{f})=\frac{1}{n_{f}}\|\mathbf{X}_{f}^{\top}\mathbf{w}-\mathbf{y}_{f}\|_{2}^{2}.\vskip-14.45377pt$
@@ 3 Naive Fine-Tuning Methods Fail To Unlearn
In empirical studies (kurmanji2024towards; warnecke2021machine; golatkar2020eternal) and Table˜1, it can be observed that fine-tuning may retain the utility of a model but struggles to forget. In this section, we revisit this phenomenon in a simplified setting, aiming to gain a basic understanding of why the vanilla fine-tuning method succeeds in retaining the model’s utility on the remaining dataset but fails to forget the targeted data it was trained on.
@@ 3.1 Distinct Features
The assumption implies that each of these datasets contains features that are unique to each dataset–there is no overlap in the features present in $\mathbf{X}_{f}$ and $\mathbf{X}_{r}$. The construction of $\mathbf{X}_{f}$ and $\mathbf{X}_{r}$ can be achieved by rearranging the samples in the matrix to group non-zero features into distinct blocks. Without loss of generality, we assume a structure where each matrix contains only one zero block to clear the analysis. It can be obtained immediately from ˜3.1 that $\mathbf{w}_{*}=\mathbf{w}_{*}^{f}+\mathbf{w}_{*}^{r}$, where $\mathbf{w}_{*}^{f}$ and $\mathbf{w}_{*}^{r}$ are the optimal solution such that $\mathbf{y}^{f}=\mathbf{X}_{f}^{\top}\mathbf{w}_{*}^{f}$ and $\mathbf{y}^{r}=\mathbf{X}_{r}^{\top}\mathbf{w}_{*}^{r}$. In an ideal scenario for classification tasks, each class possesses its own unique set of features that distinctly differentiates it from other classes. We later extended our analysis to overlapping features in Section˜3.2.
Theorem 3.2.
Suppose a model is trained by the procedure 2 and 3 separately. Under the ˜3.1, it holds that
@@
Theorem˜3.2 presents two interesting observations: 1) The fine-tuning model can perform perfectly on the remaining dataset, which indicates that the information from training data has been preserved from the original model, $\mathbf{w}_{o}$, to the unlearned model via fine-tuning, $\mathbf{w}_{t}$. 2) The loss of the fine-tuning model on the forgetting dataset consistently remains zero, which diverges from the performance of the golden model. This suggests that the fine-tuning model is unable to forget the information it previously acquired from $\mathbf{w}_{o}$, which may be contradicted by catastrophic forgetting in continual learning (parisi2019continual; ding2024understanding).
To illustrate the behavior of fine-tuning during the unlearning process more clearly, we consider the projective nature of learning. Firstly, the solution of Equation˜2 can be represented as
@@
where $\mathbf{P}_{t}$ is the projection space of $\mathbf{X}_{t}$, the $\mathbf{I}-\mathbf{P}_{t}$ is the corresponding orthogonal space, and the $\mathbf{w}_{o}$ can be also represented $\mathbf{w}_{o}=\mathbf{P}\mathbf{w}_{*}$ with $\mathbf{P}$ being the projection space of $\mathbf{X}$. According to the property of projection Corollary˜C.1, multiplying any data matrix by a projection matrix preserves the components of the data that lie within the subspace defined by the projection. Moreover, under the distinct features assumption 3.1, it holds that
@@
Therefore, the unlearned model $\mathbf{w}_{t}$ from Equation˜2 decomposed into two components for the unlearning process: the first part, $\mathbf{w}_{*}^{r}$, preserves the accuracy on the remaining data, while the second part, $\mathbf{w}_{*}^{f}$, also ensures accuracy on the forget data. However, the projection of $\mathbf{w}_{*}^{f}$ onto the fine-tuning space $\mathbf{P}_{t}$ has no effect, ultimately resulting in the unlearned model $\mathbf{w}_{t}$ being exactly the same as the pretrained model $\mathbf{w}_{o}$. The proof of Theorem˜3.2 is provided in Section˜C.2.
Figure 1: Machine Unlearning Performance via (Masked) Fine-tuning with (without) Overlapping Features. Figure˜1(a) and Figure˜1(b) present the relationship between machine unlearning loss (i.e. RA, UA) and the number of fine-tuning data samples under distinct features and overlapping features assumptions, using naive FT method. In contrast, Figure˜1(c) and Figure˜1(d) show the same relationship using masked fine-tuning methods, as discussed in Section˜4.